Journalist Report (Japanese) March 29th

ガラス越しの風景がくれたもの

29th, March 2018 MDRS time

河村信(クルージャーナリスト)

扉が閉まってから3日が経ちました。

それは、外の景色を自分の目でダイレクトに見ることが出来なくなって3日経ったということです。基地の中では窓ガラス越し、宇宙服を着れば、ヘルメット越しの景色です。

火星は、人間の生身の体が大気に触れることを許してくれない空間です。

EVA(船外活動)に出てみれば、いつも当たり前に出来ることが何ひとつとして出来ません。まず、顔を触れない。鼻水が出ても、かむことさえ出来ません。雄大なMDRSの景色に心打たれても、その表現手段は限られてしまいます。カメラのファインダーが全く見えませんし、操作もままなりません。ここに来る前に、様々な状況を自分なりにシミュレーションして機材を選んできたつもりでしたが、現実は私の想像を凌ぐ不自由さです。

電気・水・ゴミ・通信・或いは日常の細かい所作ひとつひとつが、命に係わるという状況(を作り出すこと)が、私たちが当たり前に感じている日常が日常でなくなっていくトンネルをくぐりぬけているような気持ちにします。自分の目で外を見れない3日という時間は、いま私たちは別世界に来ていると感じるに充分な時間でした。

それは同時に、私たちCrew191が、いまどこまでも無骨に、ここは火星だというシミュレーションに挑んでいることに他ならないと思います。誰ひとりとしてここを地球だと思っていません。朝起きてから夜寝るまで、皆が「本気」で「火星」の日常に挑んでいます。

だからこそ、お互いにミッションに対して活発な議論が生まれ、時には強い口調で意見を言い合います。扉が閉まっただけでこんな空間が砂漠の真ん中に出来上がるなんて、その空間の変化に自分でも驚いてしまいます。

何年後・十何年後・何十年後・人類が火星を目指し、その土地に降り立った時、いま私たちが直面している状況は必ず起こるであろうし、その現場を撮影する人は必ずいる筈です。

ジャーナリストが火星に行くとき、何が求められるのか、何を撮るのか。Crew191のメンバーとのMDRSでの生活はあと1週間余りですが、その問いに全力で向き合って、何かの答えを見つけてから、ガラス越しではない世界に戻りたいと思うようになりました。

今も窓の外を見れば、砂嵐が白いドームに吹き付けています。

少し曇ったガラスの先に見える赤い砂漠と岩山とは、私にそんな気付きを与えてくれる空間でもあったのだと思います。

Operations Report March 29th

Crew 191 Operations Report 29/03/2018

SOL:5

Name of person filing report: Wataru Okamoto

Non-nominal systems: Solar System

Notes on non-nominal systems: Solar system seems unusual in this morning. Now it looks stable.

Generator (hours run):13hours

Solar SOC 93% (Before generator is run at night)

Diesel under 25%

Propane 72psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) ? 7gallons.

Water (trailer) 60%

Water (static) 70%

Trailer to Static Pump used ? no

Water (loft) Static to Loft Pump used yes

Water Meter: 16:00 132835.5 

Toilet tank emptied: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) –

Oil Added –

ATV Fuel Used: N/A

Hours the ATVs were Used today: N/A

Notes on ATVs: N/A

Deimos rover used: –

Hours:116.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: 100%

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: –

Hours:27.9h

Beginning charge:100%

Ending charge: 100%

Currently charging: 100%

Opportunity rover used: Morisawa, Takeda

Hours: 27.4h

Beginning charge: 100%

Ending charge: 99%

Currently charging: 99%

Curiosity rover used: Tsukishiro

Hours:22.7h

Beginning charge: 100%

Ending charge: 97%

Currently charging: 97%

HabCar used and why, where? –

General notes and comments: Solar system seems unusual in this morning. We have discussion about this problem with Dr. Shannon and now it seems stable. We will keep monitoring.

Summary of internet: Under control

Summary of suits and radios: The surface of helmet polished by Plexus & microfiber clothes.

Summary of Hab operations: Shower heater doesn’t work.

Summary of GreenHab operations: No problem

Summary of ScienceDome operations: Only use checking the solar system.

Summary of RAMM operations: N/A

Summary of health and safety issues: No problem

Questions, concerns and requests to Mission Support:

The PLEXUS splay is almost empty. We need another supply.

We will put some garbage bags into the engineering airlock tonight. Please take them out.

Daily Pictures – March 29th

SOL Essay (photo description)
29 Mar 2018

Sol 05

Miho Tsukishiro (XO)

For our Journalist’s project, today we had a film-shooting in EVA. The sight from the roof of the hill along the Galileo Road was so amazing, really MARS-like scenery. But after we climbed up to there, found the difficulty to breath wearing the space helmet. Also it was hard for our video photographer to tell direction via radio because he had to use both hands to hold his camera. However, the results ware great as you can see.

By the way, we have found a problem with the solar system in HAB this morning. This is one of our lifeline, so we keep monitoring tomorrow.

In this afternoon, Wataru Okamoto has started his own project with his majoring system: the instrument has been installed inside a tunnel at last.

EVA#5 Report 29th March

EVA Report

EVA #5 – 29 Mar 2018 – Sol5

Crew members: Makoto Kawamura (leader), Miho Tsukishiro, Kai Takeda, Fumiei Morisawa (switched from Yusuke Murakami)

Objectives:

– Scouting

– Filming Mars landscape

Operations:

09:30 – Miho Tsukishiro and Kai Takeda enter the airlock

09:33 – Exit the airlock and begin to engineering check

09:54 – Makoto Kawamura and, Fumiei Morisawa enter the airlock

09:58 – EVA start (All member)

10:25 – Arrived at the Site

10:59 – Turn back to the HAB

11:25 – Enter the airlock

11:28 – EVA end

Narrative:

EVA today was a journalist EVA

We explored to describe MARS great scenery.

Using 360°camera and action cam, we challenged expressions that have never existed.

As a result, we think we are able to capture the dynamic figure of MARS.

Vehicles

– 2 Rovers (Curiosity: Makoto Kawamura, Miho Tsukishiro, Opportunity: Fumiei Morisawa, Kai Takeda)

Greenhab Report 29th March

GreenHab Report

Kai Takeda

29 MAR 2018

Environmental control: (Choose which of the following is appropriate and explain further if needed)

Ambient with window/door open

Shade cloth on

Fan on (8:30~10:30, 15:00~17:00)

Average temperatures: 24.7℃

Low 14.6℃

High 38.9℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 6gal

Time(s) of watering for crops: twice

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative:

Dear Dr.Shannon

The blue box of tomatoes is dryness. Can I put water in it?

If I can, how much water should I put in?

Support/supplies needed: None

Sol Summary 29th March

Crew 191 Sol Summary Report
29 Mar 2018

Sol 05

Summary Title:

EVA #5: Scouting + GPS Navigation training, Filming

Miho Tsukishiro (XO)

Mission Status:

After warm-up, in the process of our mission.

Sol Activity Summary:

7:30 Breakfast

8:00 Briefing

9:30 EVA #5 (leader: Makoto Kawamura, Fumiei Morisawa, Miho Tsukishiro, Kai Takeda) Habcom: Yusuke Murakami

12:00 Lunch

13:00 Personal work

18:00 Gathering reports

19:00 CapCom

Look Ahead Plan:

EVA #6: CMEs (Venzha’s project)

Anomalies in work:

N/A

Weather:

Sunny. mild winds in the morning, got strong afternoon.

Crew Physical Status:

Good.

EVA:

See EVA Report.

Reports to be file:

– SOL Summary Report

– SOL Essay

– Engineer Report

– GreenHab Report

– Photo Report

– EVA Report

– EVA Request

– Journalist Report

Support Requested:

N/A